OpenStudy (anonymous):

The brightness of the illumination of an object varies inversely as the square of the distance from the object to the source of light. How far from a bulb does an object recieve four times as much illumination as it does when it is 8 m from the bulb?

OpenStudy (anonymous):

"The brightness of the illumination of an object varies inversely as the square of the distance from the object to the source of light." OR \[I=\frac{net power}{surface area}=\frac{P}{4\pi r^2}\]

OpenStudy (anonymous):

Your looking for: \[\frac{I_2}{I_1}=...\]

OpenStudy (anonymous):

I don't know bout this -.- sorry

OpenStudy (anonymous):

Are you OK with: \[I=\frac{netpower}{surfacearea}=\frac{P}{4πr^2}\]

OpenStudy (anonymous):

No , i don't know that

OpenStudy (anonymous):

That is the formula for light intensity of a source with total power P at a distance of r.

OpenStudy (anonymous):

The distance is the only given how can i answer it??

OpenStudy (anonymous):

We don't need to know the power of the light source. The question wants us to find out how far from the bulb (r=?) do we get four times as much light intensity than at 8m from the bulb. \[I_2=4I_1\] Agreed?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

But hould i get that?

OpenStudy (anonymous):

Wait... Let math do it's magic! We know the formula for light intensity: \[I_1=\frac{P}{4 \pi r_1^2}\] \[I_2=\frac{P}{4 \pi r_2^2}\] We are investigating one light source, so P is the same for both. However, we are investigating light intensity at two different locations, one at r_1 and another at r_2. Agreed?

OpenStudy (anonymous):

Okay, and then?

OpenStudy (anonymous):

 What do you get?